OpenStudy (celeste26):

6.634 x 10 ^-23 g of argon is how many atomic mass unit (amu)?

OpenStudy (celeste26):

I got 3.9950 grams, but it was wrong ...

OpenStudy (astrophysics):

Why are you looking for grams when your question requires amu?

OpenStudy (astrophysics):

Look, \[1 amu = 1.66 \times 10^{-24}g \]

OpenStudy (astrophysics):

So you have \[\left( 6.634 \times 10 ^{-23}g \right)\left( \frac{ 1 amu }{ 1.66 \times 10^{-24}g } \right) = ?\]